OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (mertsj):

First find the dot product of the two vectors.l

OpenStudy (anonymous):

okay so the dot product is 46

OpenStudy (mertsj):

yes. Now find the magnitude of each vector.

OpenStudy (anonymous):

how do you do that?

OpenStudy (mertsj):

if the vector is <2,5> the magnitude is: \[\sqrt{4+25}=\sqrt{29}\]

OpenStudy (anonymous):

sqrt 37 and sqrt 65

OpenStudy (mertsj):

Now multiply the magnitudes.

OpenStudy (mertsj):

And then divide the dot product by the product of the magnitudes. That quotient is the cosine of the angle between the vectors so hit your inverse cosine button.

OpenStudy (anonymous):

20.3 degrees?

OpenStudy (mertsj):

That's what I got.

OpenStudy (anonymous):

Okay great thanks!

OpenStudy (mertsj):

yw